OpenStudy (anonymous):

Which graph represents the solutions to the inequality |2x - 4| greater than or equal to 4? number line with a closed circle on 0, shading to the right and a closed circle on 4, shading to the left number line with an open circle on 0, shading to the left and an open circle on 4, shading to the right number line with an open circle on 0, shading to the right and an open circle on 4, shading to the left number line with a closed circle on 0, shading to the left and a closed circle on 4, shading to the right

OpenStudy (anonymous):

It is the last choice. Since it is not a strict inequality, circles must be closed. \[ |2 x - 4|\ge 4\\ |x-2| \ge 2\\ x\le 0\\ or \\ x\ge 4 \]

OpenStudy (anonymous):

could you help me with another one?

OpenStudy (anonymous):

Post it as a different problem and tag me